The first part of the shift was dedicated to the lock to CARM_NULL started at 13:04 UTC (see report #54067). The lock lasted until 16:42 UTC and, after two following locking attempts, the planned TCS activity (carried out by Nardecchia and Taranto), started at 17:20 UTC. The measurements proceeded for the rest of the shift with the ITF unlocked without major issue.
Activity concluded at 22:05 UTC, infrared cavities left locked.
Other activities carried out during the shift:
- SGD Green Matching/ Installation of green PBS/ Telescope for AA (Delaurentis, Garaventa, Guo)
No DMS alerts were received during the shift.Sub-system reports
FFLMoni52 crashed at 17:58 UTC and 19:01 UTC. The process was restarted via VPM.